# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Appendix A Perturbation theory
In the interest of brevity, we only include $m=2$ in the above going forward, while noting that the contribution from the third eigenstate in Eq. (26) will contribute to inducing OAM in a similar way. We return to this later. At this stage, we remark that the matrix element $\bra{\psi_{2}(\boldsymbol{k})}\partial_{k_{x}}\ket{\psi_{1}(\boldsymbol{k})}$ is $\propto k_{x}$ to lowest order in $k_{x/y}$ because the equal-orbital coupling is even in centrosymmetric systems. If, however, different-parity orbitals are coupled, the lowest-order matrix element would be $\propto k_{y}$. We comment on the consequences of this shortly.
@@
$\langle L_{z}\rangle_{\psi_{1}(\boldsymbol{k},\delta t)}$ $=-k_{x}r(\boldsymbol{k})\sin\theta(\boldsymbol{k}).$ (30c)
@@ Appendix C Orbital diode effect
In the present case, we find a diode effect for the charge current whenever $t_{m}\neq 0$ and orbital Rashba coupling [6] is present. In momentum space, within our framework, the orbital Rashba coupling takes the form $H_{\text{SO}}=t_{\text{SO}}(\sin k_{y}\tilde{L}_{x}-\sin k_{x}\tilde{L}_{y})$ [38]. The parameter $t_{m}$ plays a qualitatively similar role as a spin-dependent mass in kinetic ferromagnets. In the latter case, such a spin-dependent mass provides a spin polarization and net magnetization. In the present case, $t_{m}$ does not provide any net orbital magnetization on its own, but in conjunction with an applied bias, serves to introduce a non-zero orbital texture. Together with orbital Rashba coupling, this serves to fulfill the requirements of a charge diode effect.

### v2

@@ Appendix A Perturbation theory
In the interest of brevity, we only include $m=2$ in the above going forward, while noting that the contribution from the third eigenstate in Eq. (26) will contribute to inducing OAM in a similar way. We return to this later. At this stage, we remark that the matrix element $\bra{\psi_{2}(\boldsymbol{k})}\partial_{k_{x}}\ket{\psi_{1}(\boldsymbol{k})}$ is $\propto k_{x}$ to lowest order in $k_{x/y}$. If, however, different-parity orbitals are coupled, the lowest-order matrix element would be $\propto k_{y}$. We comment on the consequences of this shortly.
@@
$\langle L_{z}\rangle_{\psi_{1}(\boldsymbol{k},\delta t)}$ $=k_{x}r(\boldsymbol{k})\sin\theta(\boldsymbol{k}).$ (30c)
@@ Appendix C Orbital diode effect
In the present case, we find a diode effect for the charge current whenever $t_{m}\neq 0$ and orbital Rashba coupling [6] is present. In momentum space, within our framework, the orbital Rashba coupling takes the form $H_{\text{SO}}=t_{\text{SO}}(\sin k_{y}\tilde{L}_{x}-\sin k_{x}\tilde{L}_{y})$ [50], as stated in the main text. The parameter $t_{m}$ plays a qualitatively similar role as a spin-dependent mass in kinetic ferromagnets. In the latter case, such a spin-dependent mass provides a spin polarization and net magnetization. In the present case, $t_{m}$ does not provide any net orbital magnetization on its own, but in conjunction with an applied bias, serves to introduce a non-zero orbital texture. Together with orbital Rashba coupling, this serves to fulfill the requirements of a charge diode effect.